(b) Randomly draw a card from a standard deck of playing cards (32 cards), let Bvent B be: the card is & Diamond Face card
Answer #1

(b) required probability=P(B)=3/52 ( since there are 3 diamond face card out of 52 cards)

(c)P(C)=521/991

(d)P(male and smoker)=(33+64+34)/991=131/991

(e)P(-0.38<Z<1.25)=P(Z<1.25)-P(Z<-0.38)=0.8944-0.3520=0.5424

(f) P(Z>a)=0.875

or, 1-P(Z<a)=0.875

or, P(Z<a)=0.125

or a=-1.15

(g) here we use standard normal variate z=(x-mean)/sd

for x=41, z=(41-45)/sqrt(5)=-1.7888

P(X>41)=P(Z>-1.7888)=1-P(Z<-1.7888)=1-0.0368=0.9632